OpenStudy (anikate):

simplify: a+b+3a^2 -3b^2

OpenStudy (asnaseer):

think back to your previous question, remember I showed you that:\[a^2-b^2=(a+b)(a-b)\]

OpenStudy (anikate):

yeah but Hw do I use it to this?

OpenStudy (asnaseer):

look at the last two terms in your equation - factor out the 3 and what will you get?

OpenStudy (anikate):

3(a^2 -b^2)

OpenStudy (asnaseer):

now can you spot anything?

OpenStudy (asnaseer):

ok @Anikate back to your question

OpenStudy (anikate):

ok :)

OpenStudy (asnaseer):

so you factored out a 3 from the last two terms to get: \(3(a^2-b^2)\) Can you now spot how to factorise this?

OpenStudy (anikate):

ummm...

OpenStudy (asnaseer):

look at my first reply to this question above ^^^

OpenStudy (anikate):

a+b+3(a+b)(a-b) I think?

OpenStudy (asnaseer):

perfect! now you see you have (a+b) at the start of this expression, and also (a+b) in the term 3(a+b)(a-b)

OpenStudy (asnaseer):

so you should be able to factor out the (a+b)

OpenStudy (anikate):

which one?

OpenStudy (asnaseer):

both

OpenStudy (anikate):

could u plz show me how? I am not so bright...

OpenStudy (asnaseer):

imagine you had: x + 3xy you could factor out the x to get: x(1+3y)

OpenStudy (asnaseer):

now just think of the 'x' as being 'a+b'

OpenStudy (anikate):

huh?

OpenStudy (asnaseer):

you have: a+b + 3(a+b)(a-b) which we can write as: (a+b) + 3(a+b)(a-b) think of the (a+b) term as being a single number, say 'x', and factor it out

OpenStudy (asnaseer):

how would you factor out the 'x' in: x + 3x(a-b)

OpenStudy (anikate):

um where did u go?

OpenStudy (asnaseer):

I was waiting for you to reply.

OpenStudy (anikate):

I was confused a bit?

OpenStudy (anikate):

3(2a+2b)(a-b)?

OpenStudy (anikate):

sry 2 keep u waiting

OpenStudy (asnaseer):

remember I cannot see you - so on sites like this it is important to give feedback via chat otherwise it is very difficult to help.

OpenStudy (anikate):

sry

OpenStudy (asnaseer):

so, my last question to you was: how would you factor out the 'x' in: x + 3x(a-b)

OpenStudy (anikate):

3(2a+2b)(a-b)

OpenStudy (asnaseer):

plz read my question carefully - your answer should contain an 'x' in it: how would you factor out the 'x' in: x + 3x(a-b)

OpenStudy (anikate):

4x(a-b)

OpenStudy (asnaseer):

ok - I see where you are confused now

OpenStudy (anikate):

ok

OpenStudy (asnaseer):

what you did was the correct factorisation of: x(a-b) + 3x(a-b)

OpenStudy (asnaseer):

what we have is: x + 3x(a-b) think of this as 'x' times one, plus 'x' times 3(a-b)

OpenStudy (anikate):

k right now on the paper I am at a+b+3(a+b)(a-b)

OpenStudy (anikate):

4x(a-b)

OpenStudy (asnaseer):

that is also wrong

OpenStudy (asnaseer):

\[a+b+3(a+b)(a-b) = (a+b)\times1+(a+b)\times3(a-b)\]

OpenStudy (asnaseer):

if we now pull out the (a+b) terms, we end up with:\[(a+b)(1+3(a-b))\]

OpenStudy (asnaseer):

does that make sense to you?

OpenStudy (anikate):

umm.. no

OpenStudy (asnaseer):